OpenStudy (anonymous):

a purse contains 20 coins consisting of nickels and dimes having a total of $1.40. how many nickels are there

OpenStudy (anonymous):

what is that?

OpenStudy (anonymous):

x = dimes and y = nickels x + y = 20 and 10x + 5y = 140 you have a system of equations now, so you can solve it

OpenStudy (anonymous):

(1) n+d=20 <=> d=20-n (n is the number of nickels and d is the number of dimes) (2)0.10*d+0.05*n=1.40 (3)0.10d+0.05(20-n)=1.40 <=>0.10d+1-0.05d=1.40 <=>0.05d=0.40 <=>d=8 =>There are 8 dimes and 12 (20-8) nickels Check: 8*0.10+12*0.05=1.40